# Cutoff for the Swendsen–Wang dynamics on the complete graph

### v1

@@ 1 Introduction
The SW dynamics is reversible with respect to $\mu$ and thus converges to it [SW87, ES88]. We refer to Step 1 as the percolation step and to Step 2 as the coloring step.
@@
In this work, we provide sharper results for the mixing time of the SW dynamics in the $\beta>\beta_{\textsc{r}}=q$ regime, further refining our understanding of the SW dynamics in the mean-field model. We show that there exists a constant $c(\beta,q)>0$ such that $\tau_{\rm mix}^{\mathrm{SW}}=c(\beta,q)\log n+\Theta(1)$, which implies that this Markov chain exhibits cutoff. A Markov chain exhibits cutoff if there is a sharp drop in the total variation distance from close to 111 to close 0 in an interval of time of smaller order than $\tau_{\rm mix}$ which is called the cutoff window; i.e., we say that there is cutoff if as $n\to\infty$, for any fixed $\varepsilon\in(0,1)$,
@@
The only previously known mixing time cutoff result for the SW dynamics was on the integer lattice graph for sufficiently small enough $\beta$ [NS19]; as such, ours is the first such result in the low-temperature SW dynamics. For the much simpler (local) Glauber dynamics, which updates a single randomly chosen vertex in each step, cutoff is known in the mean-field through the high-temperature $\beta<\beta_{\textsc{l}}$ regime [LLP10, CDL+12]. For the special case of the Ising model, cutoff has also been established for the Glauber dynamics for several graph families; see, e.g., [LS13, LS17, Kim22, vdHK+21].
@@
Recall from (2) that the mixing time of the SW dynamics is $\Theta(1)$ when $\beta<\beta_{\textsc{l}}$ and exponentially slow when $\beta\in(\beta_{\textsc{l}},\beta_{\textsc{r}})$, so whether the chain exhibits cutoff is of limited significance in those parameter regimes. It is an interesting question whether the SW dynamics exhibits cutoff at the dynamical critical points $\beta=\beta_{\textsc{l}}$ or $\beta=\beta_{\textsc{r}}$. In the former, mixing is not entirely governed by a “strong drift” towards a typical spin count distribution and relies on the variance of the dynamics as well; as such, it is unlikely that the dynamics exhibits cutoff at $\beta=\beta_{\textsc{l}}$. For $\beta=\beta_{\textsc{r}}$, the SW dynamics exhibits a first-order drift, but the percolation step is critical, which significantly complicates even the asymptotic mixing time analysis.
@@
From an algorithmic perspective, for a Markov chain like the SW dynamics that exhibits cutoff at $c(\beta,q)\log n$, knowing only the asymptotic mixing time, e.g., that $\tau_{\rm mix}^{\mathrm{SW}}=\Theta(\log n)$, is insufficient in practice, since simulating the dynamics for say $\frac{c(\beta,q)}{2}\log n$ steps, is guaranteed to result in a sample with total variation distance close to 111.
@@
We conclude this introduction with some brief remarks about our proof and techniques. We construct a refined, multi-phase coupling that converges to a pair of configurations with roughly the correct (expected) number of vertices in each spin class. (In the mean-field setting when $\beta>\beta_{\textsc{r}}$, under the stationary distribution $\mu$, there is one dominant spin and all other spins are assigned to roughly the same number of vertices.) Our coupling gradually contracts the distance to this type of configuration over two phases: first to a linear distance and then to within $O(\sqrt{n})$ distance. The second phase determines the leading order of the mixing time. Earlier analyses (e.g., [GŠV19]) used simpler arguments that yield the $O(\log n)$ mixing time bound; such arguments treat step deviations “pessimistically” and cannot provide the sharper bound we derive here. Our analysis carefully accounts for how deviations are amortized over time (i.e., we show that early fluctuations do not matter as much).
@@
An innovation of the present argument is the use of a $q\times q$ projection to a partition $\{V_{1},\dots,V_{q}\}$ of $V$ at low temperatures. This projection contains information about the number of vertices assigned each spin in each $V_{i}$ and its convergence is known to imply that of the SW dynamics due to the symmetry of the mean-field model. This idea was used before in [CDL+12, LNNP14, GŠV19] but in the high-temperature $\beta<\beta_{\textsc{l}}$ setting where the percolation steps of the SW dynamics result in sub-critical random (sub)graphs with trivial connected component structures. In the $\beta>\beta_{\textsc{r}}$ regime, we must contend with the existence of a linear-sized connected component and how it is distributed in the partition $\{V_{1},\dots,V_{q}\}$. To address this, we have an additional step in our analysis in which we argue that the fraction of vertices assigned each spin in each $V_{i}$ coalesces to the overall fraction of vertices with that spin in the full configuration before the full configuration reaches a typical spin count (i.e., before the end of second phase of the coupling). This additional guarantee allows us to design the required coupling.
@@ 2.1 Couplings
It is a standard fact that if $\Pr(T_{\mathrm{coup}}>t)\leq\varepsilon$, then $\tau_{\rm mix}(\varepsilon)\leq t$; see, e.g., [LP17]. In addition, to bound $\tau_{\rm mix}=\tau_{\rm mix}(1/4)$, it suffices to bound $\Pr(T_{\mathrm{coup}}>t)$ by any constant $\varepsilon<1$, as the bound can then be boosted by repeating the coupling a constant number of times. For establishing cutoff, however, this approach is not feasible, as it would increase the mixing time bound by a multiplicative factor. As such, throughout our proofs, we must track the probability of success of our multi-stage couplings; in particular, our overall coupling must be designed such that $\Pr(T_{\mathrm{coup}}>t)\leq 1/4$.
@@ 2.2 Random graph lemmas
Let $I$ be the number of isolated vertices in $\mathcal{G}$. For any constant $\lambda>0$, there exists a constant $C>0$ that $\Pr(I\geq Cn)\geq 1-O(n^{-1})$.
@@
$\Pr(||{\mathcal{L}}_{1}(\mathcal{G})\cap U|-\theta(\lambda)|U||\geq A\sqrt{n})\leq Ce^{-cA^{2}}.$
Finally, from Theorem 1.1 in [JL08] and the reflection principle, we obtain the following.
@@ 3 Proof of main result: cutoff when $\beta>q$
[…]
@@
see Fact 2.8. From (3) and the fact that $F(a)=a$, we know that $\theta(a\beta)$ is the unique root in (0,1](0,1]( 0 , 1 ] of the equation
@@
For any constants $\varepsilon>0$ and $\delta\in(0,1)$, for all sufficiently large $n$ and any starting state $X_{0}\in\Omega$, there exists $T=O(1)$ such that after $T$ steps, with probability at least $\delta$, the SW dynamics reaches a state $X_{T}$ such that $\|\alpha(X_{T})-m\|_{\infty}\leq\varepsilon$.
A version of this lemma was established in [GŠV19], but we require a slight refinement here so that the probability of success is any constant arbitrarily close to 111.
In the second phase, we claim that after an additional $T_{2}=c(\beta,q)\log n+O(1)$ steps, with both copies of the chain still evolving independently, for any constant $\delta\in(0,1)$ there exists a constant $C=C(\delta)>0$ such that with with probability at least $\delta$:
@@
This follows from the following lemma.
@@
A slightly weaker convergence result was proved in [GŠV19], where it is required that $T=A\log n$ with $A>c(\beta,q)$. This allowed for a technically simpler analysis. To obtain our sharper bound, we devise a refined approach in which we observe that deviations from the expected proportions vector after one step are amortized over the remaining steps of the dynamics, so that early deviations end up contributing less to the overall separation from $m$.
In preparation for the third phase of coupling, we require an additional “stationary-like” property for both copies of the chain. Let $\mathcal{V}=\{V_{1},\dots,V_{q}\}$ be a partition of $V$ where $V_{i}$ contains all vertices of $X_{T_{1}}$ assigned spin $i$. For constant $\lambda\in(0,1)$, we say that $\mathcal{V}$ is a $\lambda$-partition if $\min_{i}|V_{i}|\geq\lambda|V|$. From Lemma 3.1 we know that $\mathcal{V}$ is a $\lambda$-partition with probability at least $\delta$, for any constant $\delta$ arbitrarily close to 111.
Let $\alpha_{1,i}(X_{t})$ denote the fraction of vertices of $V_{i}$ assigned the dominant in $X_{t}$; i.e., there are $\alpha_{1,i}(X_{t})|V_{i}|$ such vertices. The following lemma ensures that during the second phase, in parallel to making progress towards $m$, at some time $T\leq T_{1}+T_{2}$, the fraction of vertices assigned the spin from the dominant spin class in each $V_{i}$ is roughly the same fraction as that from the full configuration $X_{t}$ in $V$. This additional property will allow us to couple the two process in $O(1)$ additional steps in the third phase.
@@
This lemma implies that we can obtain two configurations such that $A_{T+1}=A_{T+1}^{\prime}$ with probability $\Omega(1)$. Similar couplings were also designed in [LNNP14, GŠV19] but for the high-temperature $\beta<\beta_{\textsc{l}}$ setting where all the spin classes have roughly the same number of vertices and the percolation step is sub-critical in all classes. Here, we have to contend with the presence of a dominant spin class and a super-critical percolation step; this is the reason for the additional step in our analysis (Lemma 3.4) that guarantees that the extra assumption A2 holds after the first two phases.
The final step of our coupling is a simple boosting phase so that the overall probability of success is at least 3/4$\geq$ 3/4≥ 3 / 4.
@@ 4 Convergence to the neighborhood of the majority vector
We first prove Lemma 3.1 which shows convergence to an $O(1)$ ball around $m$ after $T=O(1)$ steps.
Proof Lemma 3.1.
From Lemma 30 in [GŠV19], we have that for any constant $\varepsilon>0$ and any starting state $X_{0}\in\Omega$, there exists $T=O(1)$ such that after $T$ steps, we have that $\|\alpha(X_{T})-m\|_{\infty}\leq\varepsilon$ with probability at least $\gamma=\Omega(1)$. Hence, running the dynamics for $kT$ steps, the probability that we do not reach a configuration $X$ such that $\|\alpha(X)-m\|_{\infty}\leq\varepsilon$ is at most $(1-\gamma)^{k}$, and the result follows. ∎
We proceed with the proof of Lemma 3.2, which ensures that SW dynamics reaches an $O(1/\sqrt{n})$ ball around $m$ after $T=c(\beta,q)\log n+O(1)$ steps. Let
@@
Let $F^{k}(x)=F^{k-1}(F(x))$ with $F^{1}=F$; i.e., $F^{k}$ is the result of applying $F$ repeatedly $k$ times. We show first that $\alpha(X_{t})$ is close in infinity norm to $\kappa(F^{t}(\alpha_{1}(X_{0})))$ for $t\geq 0$ and then that, for an appropriate choice of $t$, $\kappa(F^{t}(\alpha_{1}(X_{0})))$ is close to $m$.
@@
Suppose $X_{0}$ is such that $\|\alpha(X_{0})-m\|_{\infty}<\varepsilon$ for a sufficiently small constant $\varepsilon>0$. Then, for any $t\geq 0$ and any constant $\delta\in(0,1)$, there exists a constant $C(\delta)>0$ such that with probability at least $\delta$
@@
Suppose $x_{0}\in(a-\varepsilon,a+\varepsilon)$ where $\varepsilon>0$ is small enough and let $\gamma\in\mathbb{R}$. Then, if $T=c(\beta,q){\log n}+\gamma$, there exists a constant $C(\gamma)>0$ such that
$\frac{|x_{0}-a|C(\gamma)}{\sqrt{n}}-o\Big{(}\frac{1}{\sqrt{n}}\Big{)}\leq|F^{T}(x_{0})-a|\leq\frac{|x_{0}-a|C(\gamma)}{\sqrt{n}}+o\Big{(}\frac{1}{\sqrt{n}}\Big{)}.$
Proof of Lemma 3.2.
Follows from Lemmas 4.1 and 4.2 and the triangle inequality. ∎
We proceed with the proof of Lemma 4.1 for which we will need the following concentration of the fluctuations.
@@
Proof.
Since $\beta>q$ and $a>1/q$ by Fact 2.5, $a\beta>1$ and $b\beta<1$. By assumption $\|\alpha(X_{t})-m\|_{\infty}<\varepsilon$, so when $\varepsilon$ is small enough, exactly one color class in the percolation step of the SW dynamics is supercritical, and the remaining $q-1$ color classes are subcritical.
Let $C_{1},C_{2},\dots$ denote the connected components after the percolation step from $X_{t}$, sorted in decreasing order by size and let $R=\sum_{i\geq 2}|C_{i}|^{2}$. By Lemma 2.4 there exist constants $\gamma_{0},\gamma_{1}>0$ such that $\mathrm{E}[R]=\gamma_{0}n+O(1)$ and $\mathrm{Var}[R]\leq\gamma_{1}n.$ Then, by Chebyshev’s inequality, for any $\gamma_{2}\geq 1$ and $n$ large we have
@@
$Q_{j}^{(i)}=\begin{cases}|C_{j}|&\text{if }C_{j}\text{ colored }i,~\text{and}\\ 0&\text{otherwise}.\\ \end{cases}$
Let $Z_{i}=\sum_{j\geq 2}Q_{j}^{(i)}$ be the number of vertices colored $i$ in the coloring step. By Hoeffding’s inequality, for any $\gamma_{3}\geq 0$
@@
and combined with (6) we deduce by a union bound that
@@
Turning our attention to $|C_{1}|$, by Lemma 2.2, there are exists constants $c_{1},c_{2}>0$ such that
@@
Combining these facts and setting $\gamma_{2}=\gamma_{3}=\gamma_{4}=r/2$, it follows that for $r\geq\gamma_{0}$ and $n$ large enough that
$\Pr\Big{(}\|\alpha(X_{t+1})-\kappa(F(\alpha_{1}(X_{t})))\|_{\infty}\geq\frac{r}{\sqrt{n}}\Big{)}$ $\leq c_{1}\exp(-c_{2}\gamma_{4}^{2})+2\exp\Big{(}\frac{-\gamma_{3}^{2}}{\gamma_{0}+\gamma_{2}\sqrt{\gamma_{1}/n}}\Big{)}+\frac{4}{\gamma_{2}^{2}}$
@@
$\xi_{t}=\|\alpha_{j}(X_{t})-\kappa_{j}(F(\alpha_{1}(X_{t-1})))\|_{\infty};$
this quantity controls the fluctuation of all color classes at time $t$ from their expected value. Then,

### v2

@@ 1 Introduction
The SW dynamics is ergodic and reversible with respect to $\mu$ and thus converges to it [SW87, ES88]. We refer to Step 1 as the percolation step and to Step 2 as the coloring step.
@@
In this work, we provide sharper results for the mixing time of the mean-field SW dynamics in the $\beta>\beta_{\textsc{r}}=q$ regime, further refining our understanding of this dynamics. We show that there exists a constant $c(\beta,q)>0$ such that $\tau_{\rm mix}^{\mathrm{SW}}=c(\beta,q)\log n+\Theta(1)$, which implies that this Markov chain exhibits cutoff. A Markov chain exhibits cutoff if there is a sharp drop in the total variation distance from close to $1$ to close 0 in an interval of time of smaller order than $\tau_{\rm mix}$ which is called the cutoff window; i.e., we say that there is cutoff if as $n\to\infty$, for any fixed $\varepsilon\in(0,1)$,
@@
The only previously known mixing time cutoff result for the SW dynamics was on the integer lattice graph for sufficiently small enough $\beta$ [NS19]; ours is the first such result in the low-temperature SW dynamics. For the local Glauber dynamics, which updates a single randomly chosen vertex in each step, cutoff is known in the mean-field through the high-temperature $\beta<\beta_{\textsc{l}}$ regime [LLP10, CDL+12]. For the special case of the Ising model, cutoff has also been established for the Glauber dynamics for several graph families; see [LS13, LS17, Kim22, vdHK+21].
@@
Recall from (2) that the mixing time of the SW dynamics is $\Theta(1)$ when $\beta<\beta_{\textsc{l}}$ and exponentially slow when $\beta\in(\beta_{\textsc{l}},\beta_{\textsc{r}})$, so the question of whether the chain exhibits cutoff is of limited significance in those parameter regimes. It is an interesting open question, however, whether the SW dynamics exhibits cutoff at the dynamical critical points $\beta=\beta_{\textsc{l}}$ or $\beta=\beta_{\textsc{r}}$. In the former, mixing is not entirely governed by a “strong drift” towards a typical spin count distribution and relies on the variance of the dynamics as well; as such, it is unlikely that the dynamics exhibits cutoff at $\beta=\beta_{\textsc{l}}$. For $\beta=\beta_{\textsc{r}}$, the SW dynamics exhibits a first-order drift towards typical configurations, but the percolation step is critical, which significantly complicates even the asymptotic mixing time analysis.
@@
From an algorithmic perspective, for a Markov chain like the SW dynamics that exhibits cutoff at $c(\beta,q)\log n$, knowing only the asymptotic mixing time, e.g., that $\tau_{\rm mix}^{\mathrm{SW}}=\Theta(\log n)$, is insufficient in practice, since simulating the dynamics for say $\frac{c(\beta,q)}{2}\log n$ steps, is guaranteed to result in a sample with total variation distance close to $1$.
@@
We conclude this introduction with some brief remarks about our proof and techniques. We construct a multi-phase coupling that converges to a pair of configurations with roughly the correct (expected) number of vertices in each spin class. (In the mean-field setting when $\beta>\beta_{\textsc{r}}$, under the stationary distribution $\mu$, there is one dominant spin and all other spins are assigned to roughly the same number of vertices.) Our coupling gradually contracts the distance to this type of configuration over two phases: first to a linear distance and then to within $O(\sqrt{n})$ distance. The second phase determines the leading order of the mixing time. Earlier analyses (i.e., [GŠV19]) yield the $O(\log n)$ mixing time bound but treat step deviations “pessimistically” and cannot provide the sharper bound we derive here. Our analysis carefully accounts for how deviations are amortized over time and account for the fact that early fluctuations do not matter as much.
@@
An innovation of the present argument is the use of a $q\times q$ projection to a partition $\{V_{1},\dots,V_{q}\}$ of $V$ at low temperatures. This projection contains information about the number of vertices assigned each spin in each $V_{i}$ and its convergence is known to imply that of the SW dynamics due to the symmetry of the mean-field model. This idea was used before in [CDL+12, LNNP14, GŠV19] in the high-temperature $\beta<\beta_{\textsc{l}}$ setting where the percolation steps of the SW dynamics result in sub-critical random (sub)graphs with trivial connected component structures. In the $\beta>\beta_{\textsc{r}}$ regime, we must account for the existence of a linear-sized connected component and how it is distributed in the partition $\{V_{1},\dots,V_{q}\}$. To address this, we have an additional step in our analysis in which we argue that the fraction of vertices assigned each spin in each $V_{i}$ coalesces to the overall fraction of vertices with that spin in the full configuration before the full configuration reaches a typical spin count (i.e., before the end of second phase of the coupling). This additional guarantee allows us to design the required coupling.
@@ 2.1 Couplings
It is a standard fact that if $\Pr(T_{\mathrm{coup}}>t)\leq\varepsilon$, then $\tau_{\rm mix}(\varepsilon)\leq t$; see [LP17]. In addition, to bound $\tau_{\rm mix}=\tau_{\rm mix}(1/4)$, it suffices to bound $\Pr(T_{\mathrm{coup}}>t)$ by any constant $\varepsilon<1$, as the bound can then be boosted by repeating the coupling a constant number of times. For establishing cutoff, however, this approach is not feasible, as it would increase the mixing time bound by a multiplicative factor. Throughout our proofs, we track the probability of success of each phase of our multi-phase coupling so that $\Pr(T_{\mathrm{coup}}>t)\leq 1/4$.
@@ 2.2 Random graph lemmas
Let $I$ be the number of isolated vertices in $\mathcal{G}$. For any constant $\lambda>0$, there exists a constant $C>0$ that $\Pr(I\geq Cn)=1-O(n^{-1})$.
@@
[…]
Finally, Theorem 1.1 in [JL08] and the reflection principle yields the following lemma.
@@ 3 Proof of main result: cutoff when $\beta>q$
We assume through the remainder of the paper that $\beta>\beta_{\textsc{r}}=q$. Let $a=a(\beta,q)>1/q$ be the unique fixed point of the drift function $F$ in $(1/q,1]$; see Fact 2.5. We use in our analysis two different low-dimensional projections of the SW dynamics. The first is the $q$-dimensional proportions vector. For a configuration $X\in\Omega$, let $\alpha(X)$ be a $q$-dimensional vector, where each coordinate of $\alpha(X)$ corresponds to the fraction of vertices assigned a given spin in $X$. We will assume that $\alpha_{1}(X)$ contains the fraction of vertices assigned the dominant spin in $X$, breaking ties arbitrarily; that is, $\alpha_{1}(X)\geq\alpha_{i}(X)$ for all $i\in[q]$.
@@
see Fact 2.8. From (3), we know that $\theta(a\beta)$ is the unique root in $(0,1]$ of the equation
@@
For any constants $\delta\in(0,1)$ and $\varepsilon>0$ sufficiently small, for all sufficiently large $n$ and any starting state $X_{0}\in\Omega$, after $T=O(1)$ steps, with probability at least $\delta$, the SW dynamics reaches a state $X_{T}$ such that $\|\alpha(X_{T})-m\|_{\infty}\leq\varepsilon$.
A version of this lemma was established in [GŠV19], but we require a slight refinement here so that the probability of success is any constant arbitrarily close to $1$.
In the second phase, we claim that after an additional $T_{2}=c(\beta,q)\log n+O(1)$ steps, in which both copies of the chain continue to evolve independently, for any constant $\delta\in(0,1)$ there exists a constant $C=C(\delta)>0$ such that with with probability at least $\delta$:
@@
This is established in the following lemma.
@@
A similar convergence result was proved in [GŠV19], where it is required that $T=A\log n$ with $A>c(\beta,q)$. This allowed for a simpler analysis. To obtain our sharper bound, we use the fact that deviations from the expected proportions vector after one step are amortized over the remaining steps of the dynamics, so that early (larger) deviations end up contributing less to the overall separation from $m$.
In the analysis of the third phase of coupling, we use an additional “stationary-like” property for both copies of the chain. Let $\mathcal{V}=\{V_{1},\dots,V_{q}\}$ be a partition of $V$ where $V_{i}$ contains all vertices of $X_{T_{1}}$ assigned spin $i$. For constant $\lambda\in(0,1)$, we say that $\mathcal{V}$ is a $\lambda$-partition if $\min_{i}|V_{i}|\geq\lambda|V|$. From Lemma 3.1 we know that $\mathcal{V}$ is a $\lambda$-partition with constant probability arbitrarily close to $1$.
Let $\alpha_{1,i}(X_{t})$ denote the fraction of vertices of $V_{i}$ assigned the dominant spin in $X_{t}$; i.e., there are $\alpha_{1,i}(X_{t})|V_{i}|$ such vertices. The following lemma ensures that during the second phase, in parallel to making progress towards $m$, at some time $T_{1}+T_{2}$, the fraction of vertices assigned the spin from the dominant spin class in each $V_{i}$ is roughly the same fraction as that from the full configuration $X_{T}$ in $V$. This additional property will allow us to couple the two process in $O(1)$ additional steps in the third phase.
@@
Similar couplings were also designed in [LNNP14, GŠV19] but for the high-temperature $\beta<\beta_{\textsc{l}}$ setting where all the spin classes have roughly the same number of vertices and the percolation step is sub-critical in all classes. Here, we have to contend with the presence of a dominant spin class and a super-critical percolation step; this is the reason for the additional step in our analysis (Lemma 3.4) that guarantees that the extra assumption A2 holds after the first two phases.
The final step of our coupling is a simple boosting phase so that the overall probability of success is at least $\geq 3/4$.
@@ 4 Convergence to the neighborhood of the majority vector
We provide next the the proofs of Lemmas 3.1 and  3.2, which ensure that SW dynamics reaches an $O(1/\sqrt{n})$ ball around $m$ after $T=c(\beta,q)\log n+O(1)$ steps. For $b\in(0,1)$, let
@@
Let $F^{k}(x)=F^{k-1}(F(x))$ with $F^{1}=F$. We show first that $\alpha(X_{t})$ is close to $\kappa(F^{t}(\alpha_{1}(X_{0})))$ for $t\geq 0$ and then that, for an appropriate choice of $t$, $\kappa(F^{t}(\alpha_{1}(X_{0})))$ is close to $m$.
@@
Suppose $X_{0}$ is such that $\|\alpha(X_{0})-m\|_{\infty}\leq\varepsilon$ for a sufficiently small constant $\varepsilon>0$. Then, there exists a constant $\hat{c}(\beta,q)>c(\beta,q)$, such that for any $t\leq\hat{c}(\beta,q)\log n$ and any constant $\delta\in(0,1)$, there exists a constant $C(\delta)>0$ such that with probability at least $\delta$
@@
Suppose $x_{0}\in(a-\varepsilon,a+\varepsilon)$ where $\varepsilon>0$ is small enough and let $\gamma\in\mathbb{R}$. Then, if $T=c(\beta,q){\log n}+\gamma$, there exist constants $C_{1}=C_{1}(\beta,q,\varepsilon)>0$ and $C_{2}=C_{2}(\beta,q,\varepsilon)>0$ such that
$\frac{|x_{0}-a|F^{\prime}(a)^{\gamma}C_{1}}{\sqrt{n}}\leq|F^{T}(x_{0})-a|\leq\frac{|x_{0}-a|F^{\prime}(a)^{\gamma}C_{2}}{\sqrt{n}}.$
A key step in the proof of Lemma 4.1, which is also useful in the proof of Lemma 3.1 is the following concentration guarantee on the fluctuations.
@@
Before providing the proofs of these lemmas, we use them to prove Lemmas 3.1 and  3.2.
Proof Lemma 3.1.
From Lemma 30 in [GŠV19], we have that for any constant $\varepsilon>0$ and any starting state $X_{0}\in\Omega$, there exists $T=O(1)$ such that after $T$ steps, we have that $\|\alpha(X_{T})-m\|_{\infty}\leq\varepsilon$ with probability at least $\gamma=\Omega(1)$. Running the dynamics for $kT$ steps, ensures the probability that we do not reach a configuration $X$ such that $\|\alpha(X)-m\|_{\infty}\leq\varepsilon$ after $kT$ steps is at most $(1-\gamma)^{k}$.
To complete the proof, we we show that once we reach a configuration $\|\alpha(X_{t})-m\|_{\infty}\leq\varepsilon$, then the same is true $X_{t+1}$ with high probability provided $\varepsilon$ is small enough. The result then follows from a union bound over the steps. If $\|\alpha(X_{t})-m\|_{\infty}\leq\varepsilon$, then $\alpha_{1}(X_{t})\in[a-\varepsilon,a+\varepsilon]$. By Fact 2.5, for $\varepsilon$ small enough there exists a constant $\rho\in(0,1)$ such that $|F(a+\varepsilon)-a|\leq\rho\varepsilon$ and $|F(a-\varepsilon)-a|\leq\rho\varepsilon$. Since $F$ is increasing and concave in $(1/q,1)$ (see Fact 2.6), this implies that $(a+\varepsilon)-F(a+\varepsilon)\geq(1-\rho)\varepsilon$ and $F(a-\varepsilon)-(a-\varepsilon)\geq(1-\rho)\varepsilon$. Thus, in order for $\|\alpha(X_{t+1})-m\|_{\infty}>\varepsilon$, we would need $|\alpha_{1}(X_{t+1})-F(\alpha_{1}(X_{t}))|\geq(1-\rho)\varepsilon$, which implies that $\|\alpha(X_{t+1})-\kappa(F(\alpha_{1}(X_{t})))\|_{\infty}\geq(1-\rho)\varepsilon$. Then, from Lemma 4.3 we have that
$\Pr\Big(\|\alpha(X_{t+1})-\kappa(F(\alpha_{1}(X_{t})))\|_{\infty}\geq(1-\rho)\varepsilon\Big)\leq O\big(n^{-1}\big).$
Proof of Lemma 3.2.
Follows from Lemmas 4.1 and 4.2 and the triangle inequality. ∎
4.1 Proofs of Lemmas 4.1,  4.2, and  4.3
We provide first the proof of and Lemma 4.3.
Proof of Lemma 4.3.
Since $\beta>q$ and $a>1/q$ by Fact 2.5, $a\beta>1$ and $\frac{1-a}{q-1}\beta<1$. By assumption $\|\alpha(X_{t})-m\|_{\infty}\leq\varepsilon$, so when $\varepsilon$ is small enough, exactly one color class in the percolation step of the SW dynamics is supercritical, and the remaining q1q-1 color classes are subcritical.
Let $C_{1},C_{2},\dots$ denote the connected components after the percolation step from $X_{t}$, sorted in decreasing order by size (breaking ties arbitrarily) and let $R=\sum_{i\geq 2}|C_{i}|^{2}$. By Lemma 2.4 there exist constants $\gamma_{0},\gamma_{1}>0$, that depend only on $\beta$ and $q$, such that $\mathrm{E}[R]=\gamma_{0}n+O(1)$ and $\mathrm{Var}[R]\leq\gamma_{1}n.$ Then, by Chebyshev’s inequality, for any $\gamma_{2}\geq 1$ and $n$ large we have
@@
[…]
Let $Z_{i}=\sum_{j\geq 2}Q_{j}^{(i)}$. By Hoeffding’s inequality, for any $\gamma_{3}\geq 0$
@@
and combined with (6) we deduce that
@@
Turning our attention to $|C_{1}|$, by Lemma 2.2, there are exists constants $c_{1},c_{2}>0$ such that for any $\gamma_{4}>0$
@@
Combining these facts via a union bound, and setting $\gamma_{2}=\gamma_{3}=\gamma_{4}=r/2$, it follows that for $r\geq\gamma_{0}$ and $n$ large enough that
$\Pr\Big(\|\alpha(X_{t+1})-\kappa(F(\alpha_{1}(X_{t})))\|_{\infty}\geq\frac{r}{\sqrt{n}}\Big)$ $\leq c_{1}\exp(-c_{2}\gamma_{4}^{2})+2q\exp\Big(\frac{-\gamma_{3}^{2}}{\gamma_{0}+\gamma_{2}\sqrt{\gamma_{1}/n}}\Big)+\frac{4q}{\gamma_{2}^{2}}$
@@
$\xi_{t}=n\cdot\|\alpha(X_{t})-\kappa(F(\alpha_{1}(X_{t-1})))\|_{\infty}.$ (8)
By Fact 2.5, $|F^{\prime}(a)|<1$, so for small enough $\varepsilon$, there exists $\rho\in(0,1)$ such that $1-\rho=\max_{|x-a|<2\varepsilon}|F^{\prime}(x)|$. Let $K>0$ be a large constant we will choose later and let us define the sequence
$r_{i}=K\big(1-{\rho}/{2}\big)^{t-i}\big(1+\rho\big)^{t-i}.$ (9)
Our first observation is that when $\xi_{i}<r_{i}\sqrt{n}$ for all $i\leq t$, we have
$|F^{t-j-1}(\alpha_{1}(X_{j}))-a|<2\varepsilon$ (10)
for all $j\leq t-1$. To see this, note that since $F$ is increasing, concave and have a unique fixed point in $[1/q,1]$ (see Facts 2.5 and 2.6) we have that for any $y\in[1/q,1]$
$|F(y)-a|\leq|y-a|.$ (11)
Hence,
$|F^{t-j-1}(\alpha_{1}(X_{j}))-a|\leq|\alpha_{1}(X_{j})-a|,$
and from (8) and the triangle inequality it follows that
$|\alpha_{1}(X_{j})-a|\leq|F(\alpha_{1}(X_{j-1}))-a|+\frac{\xi_{j}}{n}.$
Using (11) again and combining these inequalities we conclude that:
$|F^{t-j-1}(\alpha_{1}(X_{j}))-a|\leq|\alpha_{1}(X_{j-1})-a|+\frac{\xi_{j}}{n}.$
Iterating this process:
$|F^{t-j-1}(\alpha_{1}(X_{j}))-a|\leq|\alpha_{1}(X_{0})-a|+\frac{1}{n}\sum_{i=1}^{j}\xi_{i}\leq\varepsilon+\frac{1}{\sqrt{n}}\sum_{i=1}^{j}r_{i}.$
Observe that for a suitable constant $K^{\prime}>0$,
$\frac{1}{\sqrt{n}}\sum_{i=1}^{j}r_{i}=\frac{K}{\sqrt{n}}\sum_{i=1}^{t}\big(1-{\rho}/{2}\big)^{t-i}\big(1+\rho\big)^{t-i}\leq\frac{K^{\prime}(\big(1-{\rho}/{2}\big)\big(1+\rho\big))^{t}}{\sqrt{n}}.$
So, when $t$ is such that $(\big(1-{\rho}/{2}\big)\big(1+\rho\big))^{t}\leq n^{\frac{1}{2}-\gamma}$ for a constant $\gamma\in(1/4,1/2)$, we have
$\frac{1}{\sqrt{n}}\sum_{i=1}^{j}r_{i}=o(1),$
and thus $|F^{t-j-1}(\alpha_{1}(X_{j}))-a|\leq\varepsilon+o(1)\leq 2\varepsilon$ for $n$ large which establishes (10). Observe that
$(\big(1-{\rho}/{2}\big)\big(1+\rho\big))^{t}\leq n^{\frac{1}{2}-\gamma}$
when $t\leq\frac{\frac{1}{2}-\gamma}{\log((1-{\rho}/{2})(1+\rho))}\log n$, and it can be checked that
$\frac{\frac{1}{2}-\gamma}{\log((1-{\rho}/{2})(1+\rho))}>c(\beta,q)=\frac{1}{2\log\frac{1}{F^{\prime}(a)}}.$
Now, going back to (8),